the reciprocal of number is 12 more than 30, find the number
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)

let the number be x
reciprocal = 1/x
1/x = 12+30
1/x= 42
x= 1/42.
...
m.ananth@hotmail.ca

Answer by ewatrrr(24785)

Hi,
Let n represent the number
.
Question States: Write as You Read
1/n = 30 + 12
1/n = 42
.
n = 1/42